Find the quadratic function that includes each set of values.
(1, 0), (2, 5), (4, 21)

Answer by stanbon(75887)

Find the quadratic function that includes each set of values.
(1, 0), (2, 5), (4, 21)
-----
Substitute each x/y pair into
ax^2 + bx + c = y
to get 3 equations in 3 unknows (a,b,c).
Solve the system to find a,b,c.
Substitute those values into y = ax^2 + bx + c
-----------------
I get y = x^2 + 2x - 3
==============
Cheers,
Stan H.
---------------